# Convex Formulation of the Zero Emission Vessel Route Planning Problem

### v2

@@ 5.1 Weight and buoyancy equilibrium
The most elementary design relation states that buoyancy and weight in calm water are in balance because the vessel is a free-floating body. This is expressed as
@@ 5.2 Cargo capacity
We model mixed-tonnage concept vessels with dedicated spaces for ro-ro cargo and passengers. In our notation, the set $\mathcal{C}$ is $\{\text{pax, roro}\}$. Here, we don’t distinguish between passenger vehicles, trucks, trailers, buses, and other types of ro-ro cargo.
@@ 5.3 Upright transverse stability
where $\tilde{h}_{\text{KB}}$ is the distance from keel to the center of buoyancy, $\tilde{h}_{\text{BM}}$ is the distance from the center of buoyancy to the metacenter and $\tilde{h}_{\text{KG}}$ is the distance from keel to the center of gravity (figure 5). The vessel is stable when $\tilde{h}_{\text{GM}}>0$ by design, meaning the vessel reverts to an upright orientation from a small angle of heel.
Figure 5: Shifting of center of buoyancy due to heel. The lever arm GZ between buoyancy and weight vectors must be positive by design. Under this condition, the vessel reverts to upright orientation.
@@ 5.4 Structural strength
The structural strength model ensures that the hull withstands external bending and shear loads without failure. Global structural behavior is the response of the hull to loads acting at the entire length of the hull. At this level, the vessel is modeled as a beam, called hull girder, that is loaded with deadweight and lightweight and supported by buoyancy. The properties of the midship cross-section determine the strength of the hull girder.
Figure 6 depicts the hull girder cross-section topology of a ro-pax vessel. The girder extends from the keel to the topmost continuous deck, excluding the superstructure, which does not span the entire length of the hull. Only continuous longitudinal elements of the hull contribute to the structural strength. These include the bottom, inner bottom, side shell, longitudinal bulkheads, and the ro-ro-decks. The elements and their relative positions are fixed, but principal dimensions and plate thicknesses are decision variables whose values determine the strength properties.
An imbalance of buoyancy and weight along the length of the hull creates a bending moment that rises to the maximum value at amidships. We integrate twice the net force per unit length (estimated in A), which results in the approximate still water bending moment at amidships:
@@
A vessel on waves experiences slowly varying buoyancy forces. Classification rules give the formula for the maximum wave bending moment according to the geometry of the hull:
@@
where $\sigma^{\text{perm}}$ is the permissible normal stress. The quantity $Z^{\text{deck}}$ is a monomial expression for the deck section modulus, which relates the applied moment to the maximum bending stress. It depends on the midship cross-section and is derived in B.
@@
where the value of the coefficient $\varphi_{V}\in[-0.92,1]$ varies along the length of the hull. Classification rules give it for both positive and negative shear forces. By taking the still water shear force, and adding to it the contribution from waves, we find the net design shear force $V^{\text{des}}=\lvert V^{\text{sw}}\lvert+\lvert V^{\text{wv}}\rvert$ located at $y=0.3L$. The corresponding stress limit constraint is
@@
Here $\tau^{\text{perm}}$ is the permissible shear stress. The vertical shear force is mapped to the maximum shear stress response via the quantity $q_{\text{NA}}^{\text{sp}}$, which is a monomial expression for the shear flow at the neutral axis of the side plating. It is derived in C.
@@
5.5 Frictional resistance
The total vessel resistance $R$ is modeled as a sum of contributions from two sources: frictional resistance $R_{F}$ and residual resistance $R_{R}$. We derive the former component in this section and the latter in §5.6. The total resistance enters into the effective towing power formula in §5.7, which determines the power drawn from the battery along with propeller and power conversion efficiency.
@@ 5.5 Frictional resistance
A numerical integration scheme for calculating the wetted area $A_{S}$ is described in D.
5.6 Residual resistance
The residual resistance is given by (note $BT$ as the reference area instead of the wetted surface):
@@ 5.6 Residual resistance
The empirical method by Hollenbach (1998) gives the coefficient $C_{R}$ for a vessel floating on even keel as
@@
Since the sum $\omega_{2}+C_{B}\omega_{5}+C_{B}^{2}\omega_{8}$ is negative according to Hollenbach (1998), (22) cannot be expressed as a log-convex constraint directly. However, we can fit a softmax-affine function (5) to log-transformed data generated with (22) in the relevant $\mathrm{Fr}$ range. A fitting function with two exponential terms achieves a sufficiently good fit. The root mean square error is less than 0.1%.
@@
5.7 Cell and battery models
@@
The battery capacity must be sufficient to supply the energy demand of the sailing leg with the greatest energy consumption. The required energy supply is multiplied by the excess capacity factor $\Theta$ and the inverse of discharging efficiency $\eta^{\text{dis}}$ to impose a constraint on the minimum capacity $Q$ as
$Q\geq\frac{\Theta}{\eta^{\text{dis}}}\max\left\{\max\left\{(bt^{\text{sea}})_{ij},(bt^{\text{sea}})_{ji}\right\}:(i,j)\in\mathcal{L}\right\}.$ (26)
@@ 5.8 Battery capacity degradation
$\phi(t)=\left[\gamma^{\text{acc}}(t)\right]^{\chi_{1}}\left[\chi_{2}+\gamma^{\text{cha}}(t)\chi_{3}\right]\exp\left(\frac{-E_{A}+\chi_{4}\gamma^{\text{cur}}(t)}{R_{G}T^{c}}\right).$ (27)
@@
$\xi=\left(\chi_{2}+\frac{1}{2}\chi_{3}\right)\exp\left(\frac{-E_{A}}{R_{G}T^{\text{c}}}\right),\quad r^{\text{dis}}=\sum_{(i,j)\in\mathcal{L}}(\tilde{b}_{ij}t_{ij}^{\text{sea}}+\tilde{b}_{ji}t_{ji}^{\text{sea}}).$
@@
We let $N^{\text{batt}}$ denote the number of complete batteries we need over the vessel’s lifetime. Since one battery is always installed for the newbuilding, $N^{\text{batt}}$ is defined as
@@
where $N^{\text{ph}}$ is the number of planning horizon periods in a year and $t^{\text{life}}$ is vessel lifetime in years. The variable $N^{\text{life}}$ is used in the cost objective (40) to model the battery investment capital cost.
@@ 5.9 Battery deck arrangement
The kth room is specified by its width $w_{k}^{\text{batt}}$, length $l_{k}^{\text{batt}}$, equal height $h^{\text{batt}}$, and lower corner $(w_{k}/2,\tilde{l}_{k},\tilde{h})^{\text{batt}}$ in relation to the reference point located at the keel centerline of the forward perpendicular. We impose lower bounds on the room height and elevation from keel:
@@
$\frac{w_{k}^{\text{batt}}}{2}$ $\leq H^{\text{fore}}(\tilde{l}_{k}^{\text{batt}},\tilde{h}^{\text{batt}}),\quad k=1,\ldots,\lceil N^{\text{room}}/2\rceil,$ (32)
@@
A vessel is designed to float on an even keel by longitudinally aligning the vessel weight and the buoyant force vectors. The position of the centermost room determines the battery center of gravity because the room arrangement is symmetrical. The resulting longitudinal stability constraint is
@@
The first two constraints state that the first ro-ro deck is above the battery deck and the waterline. The third constraint specifies vessel depth as the distance from keel to the second ro-ro deck, which we take as the uppermost continuous deck.
@@ 5.10 Weight breakdown
$+\underbrace{\rho^{\text{mass}}Q}_{\text{Battery}}+\underbrace{\rho^{\text{mot}}P^{\text{shaft, max}}}_{\text{Motors}},$ (38)
$W_{D}$ $=\underbrace{1565f^{\text{cap}}_{c=\text{roro}}}_{\text{Ro-ro cargo}}+\underbrace{170(0.02+1.146f^{\text{cap}}_{c=\text{pax}})}_{\text{Passengers and crew}}$
$+\underbrace{170(0.02+1.146f^{\text{cap}}_{c=\text{pax}})}_{\text{Freshwater}}.$ (39)
@@
$A_{k}^{\text{tbh}}=\int_{0}^{\tilde{h}^{\text{roro}}}H^{\text{fore}}(\tilde{l}_{k}^{\text{batt}},z)\,dz=\frac{\beta}{\beta+1}\tilde{h}^{\text{roro}}_{k}B\left(\frac{\tilde{h}^{\text{roro}}_{k}}{T}\right)^{1/{\beta}}\sqrt{\frac{2\tilde{l}^{\text{batt}}_{k}}{L}}.$
@@
6 Problem formulation
The ZEVRPP is the problem of minimizing the total fleet cost over a set defined by a system of inequalities from §4 and §5:
minimize (40)
@@ 6 Problem formulation
[…]
Table 2 enumerates the decision variables. All the other quantities introduced in §5 are functions of these variables.
The objective function is a sum of annualized investment and operation costs:
$f^{\text{obj}}$ $=\sum_{s\in\mathcal{S}}N^{v}_{s}\Biggl[\frac{1}{t^{\text{life}}}\Bigl(\underbrace{C^{\text{hotel}}Q_{s}^{\text{pax}}}_{\text{Hotel crew}}+\underbrace{C^{\text{deck}}}_{\text{Deck crew}}+\underbrace{C^{\text{batt}}N_{s}^{\text{batt}}k_{s}^{\text{batt}}E_{s}^{\text{batt}}}_{\text{Battery}}$ (40)
$+\underbrace{C^{\text{steel}}W_{s}^{\text{plate}}}_{\text{Hull and outfitting}}\Bigr)+\sum_{(i,j)\in\mathcal{L}_{s}}N^{\text{ph}}N^{t}_{s}\Bigl(\underbrace{C_{j}^{\text{port}}V^{\text{GT}}_{s}}_{\text{Berthing charge}}+\underbrace{C_{j}^{\text{el}}E^{\text{port}}_{ijs}}_{\text{Electricity}}\Bigr)\Biggr]$
$+\frac{1}{t^{\text{life}}}\sum_{i^{\prime}\in\mathcal{N}}\underbrace{C^{\text{cha}}P_{i^{\prime}}^{\text{cha}}}_{\text{Chargers}}.$
The port charges are levied in proportion to the vessel’s gross tonnage $V^{\text{GT}}$ calculated as
[…]
where $V^{\text{int}}$ is the internal volume of the vessel:
@@
(Here $C^{\text{int}}_{B}$ is an internal volume block coefficient). The last term in (41) replaces the log-function with its local monomial approximation at $\hat{V}^{\text{int}}$. A reasonable choice of $\hat{V}^{\text{int}}$ is $10^{5}$. The error from the approximation is less than 0.1% for typical vessel sizes.
Table 2: Decision variables. We use $s$ to index both services and vessels because all the vessels deployed on a service are identical.
Symbol Description
$f_{ijcs}$ Quantity of cargo $c$ transported from port $i$ to
port $j$ by service $s$ during the planning horizon
$v_{ijs}$ Sailing speed in leg $(i,j)$ in service $s$
$t_{ijs}^{\text{port}}$ Turnaround time at the destination port
of sailing leg $(i,j)$ in service $s$
$L_{s}$ Hull length of vessel $s$
$L^{\text{sup}}_{s}$ Superstructure length of vessel $s$
$B_{s}$ Hull waterline breadth of vessel $s$
$T_{s}$ Draft of vessel $s$
$D_{s}$ Depth of vessel $s$
$p_{s}$ Deck plate thickness of vessel $s$
$h^{\text{batt}}_{s}$ Battery room height in vessel $s$
$w^{\text{batt}}_{ks}$ Width of battery room $k$ in vessel $s$
$l^{\text{batt}}_{ks}$ Length of battery room $k$ in vessel $s$
$\tilde{h}^{\text{batt}}_{s}$ Battery room floor elevation from keel in vessel $s$
$\tilde{l}^{\text{batt}}_{ks}$ Battery room $k$ front wall distance
from forward perpendicular in vessel $s$
$\tilde{h}^{\text{roro}}_{s}$ First ro-ro deck elevation from keel in vessel $s$
$E_{s}^{\text{batt}}$ Battery capacity of vessel $s$
$t_{s}^{\text{cell}}$ Cell cycle life in vessel $s$
$N_{s}^{\text{batt}}$ Number of battery replacements in vessel $s$
$P_{i}^{\text{cha}}$ Charger power at port $i$
$r_{ijs}^{\text{Cf}}$ Auxiliary variable for modeling frictional
resistance in sailing leg $(i,j)$ in service $s$
$r_{ijs}^{\text{Cr,std}}$ Auxiliary variable for modeling residual
resistance in sailing leg $(i,j)$ in service $s$
integration of hull wetted area of vessel $s$
integration of hull steel area of vessel $s$
$N^{\text{rt}}_{s}$ Frequency of service $s$
$N^{\text{vessel}}_{s}$ Number of vessels deployed in service $s$
7 Numerical examples

### v3

@@ 5.1 Weight and buoyancy equilibrium
The most elementary design relation states that buoyancy and weight in calm water are in balance because the ship is a free-floating body. This is expressed as
@@ 5.2 Cargo capacity
We model mixed-tonnage concept ships with dedicated spaces for ro-ro cargo and passengers. In our notation, the set $\mathcal{C}$ is $\{\text{pax, roro}\}$. Here, we don’t distinguish between passenger vehicles, trucks, trailers, buses, and other types of ro-ro cargo.
@@ 5.3 Upright transverse stability
where $\tilde{h}_{\text{KB}}$ is the distance from keel to the center of buoyancy, $\tilde{h}_{\text{BM}}$ is the distance from the center of buoyancy to the metacenter and $\tilde{h}_{\text{KG}}$ is the distance from keel to the center of gravity (figure 5). The ship is stable when $\tilde{h}_{\text{GM}}>0$ by design, meaning the ship reverts to an upright orientation from a small angle of heel.
Figure 5: Shifting of the center of buoyancy due to heel. The lever arm GZ between buoyancy and weight vectors must be positive by design. Under this condition, the ship reverts to upright orientation.
@@ 5.4 Structural strength
The structural strength model ensures that the hull withstands external bending and shear loads without failure. Global structural behavior is the response of the hull to loads acting at the entire length of the hull. At this level, the ship is modeled as a beam, called hull girder, that is loaded with deadweight and lightweight and supported by buoyancy. The properties of the midship cross-section determine the strength of the hull girder.
Figure 6 depicts the hull girder cross-section topology of a ro-pax ferry. The girder extends from the keel to the topmost continuous deck, excluding the superstructure, which does not span the entire length of the hull. Only continuous longitudinal elements of the hull contribute to the structural strength. These include the bottom, inner bottom, side shell, longitudinal bulkheads, and the ro-ro decks. The elements and their relative positions are fixed, but principal dimensions and plate thicknesses are decision variables whose values determine the strength properties.
An imbalance of buoyancy and weight along the length of the hull creates a bending moment that rises to the maximum value at amidships. We integrate twice the net force per unit length (estimated in B), which results in the approximate still water bending moment at amidships:
@@
A ship on waves experiences slowly varying buoyancy forces. Classification rules give the formula for the maximum wave bending moment according to the geometry of the hull:
@@
where $\sigma^{\text{perm}}$ is the permissible normal stress. The quantity $Z^{\text{deck}}$ is a monomial expression for the deck section modulus, which relates the applied moment to the maximum bending stress. It depends on the midship cross-section and is derived in C.
@@
where the value of the coefficient $\varphi_{V}\in[-0.92,1]$ varies along the length of the hull. Classification rules are given for both positive and negative shear forces. By taking the still water shear force, and adding to it the contribution from waves, we find the net design shear force $V^{\text{des}}=\lvert V^{\text{sw}}\lvert+\lvert V^{\text{wv}}\rvert$ located at $y=0.3L$. The corresponding stress limit constraint is
@@
Here $\tau^{\text{perm}}$ is the permissible shear stress. The vertical shear force is mapped to the maximum shear stress response via the quantity $q_{\text{NA}}^{\text{sp}}$, which is a monomial expression for the shear flow at the neutral axis of the side plating. It is derived in D.
@@
5.5 Resistance
The total ship resistance $R$ is modeled as a sum of contributions from two sources: frictional resistance $R_{F}$ and residual resistance $R_{R}$. We derive both components in the following. The total resistance enters into the effective towing power formula in §5.6, which determines the power drawn from the battery along with propeller and power conversion efficiency.
@@ 5.5 Resistance
A numerical integration scheme for calculating the wetted area $A_{S}$ is described in E. The residual resistance is given by (note $BT$ as the reference area instead of the wetted surface):
@@
The empirical method by Hollenbach (1998) gives the coefficient $C_{R}$ for a ship floating on even keel as
@@
Since the sum $\omega_{2}+C_{B}\omega_{5}+C_{B}^{2}\omega_{8}$ is negative according to Hollenbach (1998), (32) cannot be expressed as a log-convex constraint directly. However, we can fit a softmax-affine function (Hoburg et al., 2016) to log-transformed data generated with (32) in the relevant $\mathrm{Fr}$ range. A fitting function with two exponential terms achieves a sufficiently good fit. The root mean square error is less than 0.1%.
@@
5.6 Cell and battery pack
@@
The battery capacity must be sufficient to supply the energy demand of the sailing leg with the greatest energy consumption. The required energy supply is multiplied by the excess capacity factor $\Theta$ to impose a constraint on the minimum capacity $Q$ as
$Q\geq\frac{\Theta}{\eta^{\text{dis}}}\max\left\{E_{ij}^{\text{leg}}:(i,j)\in\mathcal{L}_{s}\right\},$ (36)
where $E_{ij}^{\text{leg}}=(bt^{\text{transit}})_{ij}/\eta^{\text{dis}}$ is the electrical energy drawn from the battery on the sailing leg $(i,j)$.
@@ 5.7 Battery capacity degradation
$\phi(t)=\\ \left[\gamma^{\text{acc}}(t)\right]^{\chi_{1}}\left[\chi_{2}+\gamma^{\text{cha}}(t)\chi_{3}\right]\exp\left(\frac{-E_{A}+\chi_{4}\gamma^{\text{cur}}(t)}{R_{G}T^{c}}\right).$ (37)
@@
$\xi$ $=\left(\chi_{2}+\frac{1}{2}\chi_{3}\right)\exp\left(\frac{-E_{A}}{R_{G}T^{\text{c}}}\right),$
$r^{\text{dis}}$ $=\sum_{(i,j)\in\mathcal{L}_{s}}\tilde{b}_{ij}t_{ij}^{\text{transit}}.$
@@
We let $N^{\text{batt}}$ denote the number of complete batteries we need over the ship’s lifetime. Since one battery is always installed for the newbuilding, $N^{\text{batt}}$ is defined as
@@
where $N^{\text{ph}}$ is the number of planning horizon periods in a year and $t^{\text{life}}$ is the ship lifetime in years. The variable $N^{\text{life}}$ is used in the cost objective (50) to model the battery investment capital cost.
@@ 5.8 Battery deck arrangement
The kth room is specified by its width $w_{k}^{\text{batt}}$, length $l_{k}^{\text{batt}}$, equal height $h^{\text{batt}}$, and lower corner $(w_{k}/2,\tilde{l}_{k},\tilde{h})^{\text{batt}}$ in relation to the reference point located at the keel centerline of the forward perpendicular. We impose lower bounds on the room height and elevation from the keel:
@@
$\frac{w_{k}^{\text{batt}}}{2}\leq H^{\text{fore}}(\tilde{l}_{k}^{\text{batt}},\tilde{h}^{\text{batt}}),\,k=1,\ldots,\lceil N^{\text{room}}/2\rceil,$ (42)
@@
A ship is designed to float on an even keel by longitudinally aligning the ship’s weight and the buoyant force vectors. The position of the centermost room determines the battery center of gravity because the room arrangement is symmetrical. The resulting longitudinal stability constraint is
@@
The first two constraints state that the first ro-ro deck is above the battery deck and the waterline. The third constraint specifies ship depth as the distance from keel to the second ro-ro deck, which we take as the uppermost continuous deck.
@@ 5.9 Weight breakdown
$+\underbrace{\rho^{\text{mass}}Q/\rho^{\text{vol}}}_{\text{Battery}}+\underbrace{\rho^{\text{mot}}P^{\text{shaft, max}}}_{\text{Motors}},$ (48)
$W_{D}$ $=\underbrace{1565q^{\text{cap}}_{c=\text{roro}}}_{\text{Ro-ro cargo}}+\underbrace{170(0.02+1.146q^{\text{cap}}_{c=\text{pax}})}_{\text{Passengers and crew}}$
$+\underbrace{170(0.02+1.146q^{\text{cap}}_{c=\text{pax}})}_{\text{Freshwater}}.$ (49)
@@
$A_{k}^{\text{tbh}}$ $=\int_{0}^{\tilde{h}^{\text{roro}}}H^{\text{fore}}(\tilde{l}_{k}^{\text{batt}},z)\,dz$
$=\frac{\beta}{\beta+1}\tilde{h}^{\text{roro}}_{k}B\left(\frac{\tilde{h}^{\text{roro}}_{k}}{T}\right)^{1/{\beta}}\sqrt{\frac{2\tilde{l}^{\text{batt}}_{k}}{L}}.$
@@
5.10 Cost terms
The total annualized ship acquisition and variable operation cost (excluding energy) in (1) is
$C^{\text{ship}}$ $=C^{\text{hotel}}f^{\text{cap}}_{\text{pax}}+C^{\text{deck}}+C^{\text{batt}}N^{\text{batt}}Q+C^{\text{hull}}V^{\text{GT}}.$ (50)
The cost terms are hotel crew, deck crew, battery, and construction.
The construction cost and port charges are levied in proportion to the ship’s gross tonnage $V^{\text{GT}}$ calculated as
@@ 5.10 Cost terms
where $V^{\text{int}}$ is the internal volume of the ship:
@@
(Here $C^{\text{int}}_{B}$ is an internal volume block coefficient.) The last term in (51) replaces the log-function with its local monomial approximation at $\hat{V}^{\text{int}}$. A reasonable choice of $\hat{V}^{\text{int}}$ is $10^{5}$. The error from the approximation is less than 0.1% for typical ship sizes.
6 Computational experiments
This section focuses on benchmarking the nonlinear formulation of the ESFSMP (with continuous ship sizing) against the conventional linear formulation (with fixed ship types). The aim is to assess the practical value of developing and employing the posynomial models that enable using a physics-based continuous ship sizing in ESFSMP. We formulate test cases of varying size and solve both nonlinear and linear formulations using the same open-source and commercial solvers for convex mixed-integer nonlinear programming problems. The nonlinear and linear problem formulations are compared in terms of the resulting problem sizes, i.e, the number of decision variables and constraints. Computational performance is compared in terms of total cost objective value and solution time.
6.1 Input data
We construct a shipping network consisting of eight ports ordered on a straight line. The distances between the ports are randomly generated from a uniform distribution on the interval $[40,200]$ (km). The demands of ro-ro and passenger traffic are also randomly generated from a uniform distribution on the interval $[0.5,5]$ (1000 passengers or 1000 lane-meters). A scenario labeled small consists of two pendulum services that use the first four ports. Two other scenarios labeled medium and large consist of four services over the first six ports and six services over all the ports, respectively. The first service in each scenario visits every port. The other services are randomly generated. Figure 9 illustrates the ports and the outbound sailing legs of each service.
Figure 9: Randomly generated shipping networks. The directed arcs of different styles denote the outbound sailing legs of each service.
6.2 Linear benchmark problem
F formulates the linear ESFSMP. The modeling principle, based on Havre et al. (2022), is to define a small number of discrete speed and energy levels and linearly interpolate the other values. Each discrete level is associated with a weight variable. The linear model also introduces binary variables for selecting ship types from a predetermined set of alternatives.
There are two additional minor restrictions compared to the ESFSMP formulated in 4.5. First, every port is equipped with the same fixed charger power. Second, cargo offloading and loading time at each port are independent of the cargo volume. We apply these additional restrictions to the nonlinear problem instances as well to ensure that the performance comparisons are valid.
6.3 Results
We formulate linear and nonlinear problem instances for the three input data scenarios described above. The implementations are available via the link provided in §1. The linear instances use 21 ship types, which are generated using the ship model from §5. Table 4 lists the number of integer variables, continuous variables, and constraints in each problem instance. The modeling software transforms the nonlinear (log-convex) instances to a standard power cone form, which adds a large number of continuous auxiliary variables. The linear instances contain at least 20x the number of integer variables compared to the nonlinear instances.
We solve all the problem instances using MOSEK 11.0, which is a high-performance commercial solver for convex and mixed-integer optimization problems. We also use ECOS 2.0 (Domahidi et al., 2013) to solve the nonlinear instances and SCIP 10.0 (Hojny et al., 2025) to solve the linear instances. Both are free to use non-commercial solvers. The computing platform is a single-core 2.25 GHz processor with 13.3 GB of memory.
Table 5 reports the solution times and the optimal objective function values for the linear problem instances. MOSEK solves the small instances almost 3x faster than SCIP, but is more than 1.6x slower solving the medium instance and 2.4x slower solving the large instance. This reflects the typical performance variability of mixed-integer optimizers. The main message from table 5 is that the problem quickly becomes intractable when the number of integer variables increases. The worst-case solution time is less than one and a half minutes for the small instance but over three hours for the large instance.
Table 6 reports the results for the nonlinear problem instances. Total cost objective value improves by 11.7 to 19.7%, while the worst-case solve time is reduced by 10.5 to 19.6x compared to the linear instances. The non-commercial solver ECOS is capable of solving all the instances and shows worse performance than MOSEK only for the large problem instance.
When interpreting the results, it is important to recognize that the solvers find the global optimum (within a small tolerance) of all the instances. The differences in solution quality between the linear and nonlinear instances do not originate from poor local solutions. Instead, the differences are attributed to the fact that the nonlinear formulation can find improved ship designs outside of the 21 predetermined types in the linear formulation.
Table 4: Problem size as reported by MOSEK after the pre-solve stage.
Problem Formulation Integer variables Continuous variables Constraints
Small Linear 477 830 1079
Small Nonlinear 20 1544 612
Medium Linear 881 1775 2691
Medium Nonlinear 40 4675 1831
Large Linear 1410 3328 5023
Large Nonlinear 56 7900 3058
Table 5: Solver performance for the linear problem instances.
Problem Solver Objective Solving
[MEUR] time [s]
Small MOSEK 68.2 22.6
Small SCIP 68.2 81.4
Medium MOSEK 98.2 2181.3
Medium SCIP 98.2 1317.2
Large MOSEK 227.2 10804.3
Large SCIP 227.2 4495.7
Table 6: Solver performance for the nonlinear problem instances.
Problem Solver Objective Solving
[MEUR] time [s]
Small ECOS 54.8 1.6
Small MOSEK 54.8 7.7
Medium ECOS 86.8 44.4
Medium MOSEK 86.8 129.3
Large ECOS 198.3 551.2
Large MOSEK 198.3 223.5
7 Baltic Sea ferry service design